OpenStudy (liizzyliizz):

derivative of y=(x^2-x+2)^2/ x

OpenStudy (across):

\[f(x)=\frac{(x^2-x+2)^2}{x}.\]Well, what have you tried doing thus far?

OpenStudy (anonymous):

you can use the quotient rule here have you heard of that?

OpenStudy (across):

If you're lazy, you can also try rearranging things a little:\[f(x)=(x^2-x+2)^2x^{-1}.\]This turns into the form:\[f(x)=g(x)h(x),\]where you can perform the product rule:\[f'(x)=g'(x)h(x)+g(x)h'(x).\]Same thing, nonetheless.

OpenStudy (liizzyliizz):

yeah i was using it but i had a question for the first part. since theres an exponent on the outside do u also find the internal derivative?

myininaya (myininaya):

you mean do you apply chain rule yes you apply chain rule after applying quotient/product rule

OpenStudy (liizzyliizz):

ok so how would u show and continue ur work with the chain rule. i feel my algebra skills .

OpenStudy (liizzyliizz):

are failing me *

OpenStudy (across):

\[f(x)=g(h(x)),\]\[f'(x)=g'(h(x))h'(x).\]That is the chain rule.

myininaya (myininaya):

\[f(x)=(x^2-x+2)^2x^{-1}\] first i will apply product rule \[f'(x)=[(x^2-x+2)^2]'x^{-1}+(x^2-x+2)^2(x^{-1})'\] \[=2(x^2-x+2)(2x-1+0)x^{-1}+(x^2-x+2)^2(-1x^{-1-1})\]

myininaya (myininaya):

then i used chain rule and power rule and constant rule

OpenStudy (liizzyliizz):

yeah i stopped after this part , i didnt know where to continue from here

myininaya (myininaya):

well we found the derivative if you want we can make this look nicer

OpenStudy (liizzyliizz):

sure

myininaya (myininaya):

\[=\frac{2(x^2-x+2)(2x-1)}{x}-\frac{(x^2-x+2)^2}{x^2}\] how do you feel about this? if you feel good about this step then i will proceed to combine fractions

myininaya (myininaya):

or maybe i will tell you how to do the next step instead of do it for you

OpenStudy (across):

Don't forget:\[\frac{a}{b}\pm\frac{c}{d}=\frac{ad\pm bc}{bd}.\]

OpenStudy (liizzyliizz):

yeah that i understand, its the (2x-1)(x^-1) that throws me off, like i know its supposed to be there, but what gets multipled by that... that is my concern

myininaya (myininaya):

are you talking about the step where i was finding the derivative or you having a problem with the process of "simplification"

myininaya (myininaya):

are you okay with chain rule?

myininaya (myininaya):

here i will use quotient rule instead of product rule....

myininaya (myininaya):

\[f'(x)=\frac{[(x^2-x+2)^2]'x-(x^2-x+2)^2(x)'}{x^2}\]

myininaya (myininaya):

(x)'=1 right?

OpenStudy (liizzyliizz):

simplification

myininaya (myininaya):

\[[(x^2-x+2)^2]'=2(x^2-x+2)^{2-1}(2x-1+0)=2(x^2-x+2)^1(2x-1)\] \[=2(x^2-x+2)(2x-1)\]

myininaya (myininaya):

so we have \[f'(x)=\frac{2(x^2-x+2)(2x-1)x-(x^2-x+2)^2}{x^2}\]

myininaya (myininaya):

now we already have our fractions combined in the numerator you have two terms what do the two terms have in common?

OpenStudy (liizzyliizz):

yeah

myininaya (myininaya):

what do they have in common?

myininaya (myininaya):

what do they have in common?

myininaya (myininaya):

x^2-x+2

myininaya (myininaya):

what do they have in common?

myininaya (myininaya):

\[f'(x)=\frac{(x^2-x+2)[2(2x-1)x-(x^2-x+2)]}{x^2}\]

myininaya (myininaya):

what do they have in common?

myininaya (myininaya):

but 2(2x-1)x-(x^2-x+2) =2x(2x)-2x(1)-x^2+x-2 =4x^2-2x-x^2+x-2 =3x^2-x-2 \[f'(x)=\frac{(x^2-x+2)(3x^2-x-2)}{x^2}\]

OpenStudy (liizzyliizz):

sorry my computer is lagging, they have the x^2-x+2 in common... and ohh thats what happens. ahhh i wish this was cleaer to me in my head, i wouldnt be strugling as much. its the algebra thats getting to me... :/

OpenStudy (liizzyliizz):

clearer *